# A Modular Approach for Multimodal Summarization of TV Shows

### v4

@@ Abstract
In this paper we address the task of summarizing television shows, which touches key areas in AI research: complex reasoning, multiple modalities, and long narratives. We present a modular approach where separate components perform specialized sub-tasks which we argue affords greater flexibility compared to end-to-end methods. Our modules involve detecting scene boundaries, reordering scenes so as to minimize the number of cuts between different events, converting visual information to text, summarizing the dialogue in each scene, and fusing the scene summaries into a final summary for the entire episode. We also present a new metric, PRisma (Precision and Recall Evaluation of Summary Facts), to measure both precision and recall of generated summaries, which we decompose into atomic facts. Tested on the recently released SummScreen3D dataset Papalampidi and Lapata (2023), our method produces higher quality summaries than comparison models, as measured with ROUGE and our new fact-based metric.
@@ 3.2 Scene Detection
We define a cost for a given partition, by invoking the minimum description length (MDL) principle Grünwald (2007), in which the optimal representation of a piece of data is that which contains the fewest of bits. MDL has been shown to have many nice theoretical properties (Grünwald, 2007) and has been used in recent machine learning applications from interpreting hidden representations Voita and Titov (2020) to measuring noise-free information content in images (Mahon and Lukasiewicz, 2024). We use it here to formalize the desirable properties of a good segmentation.
@@ 3.5 Summary Generation
Because of our two-layer summarization approach, we can use a relatively small backbone model, and are still able to encode very long input. In our experiments, we use variants of BART-large Lewis et al. (2020) for both the dialogue summarization and high-level summarization modules, but any other similar model could be used instead. For the dialogue summarization module (Figure 1, row 4), we use the public Huggingface checkpoint for BART which has been fine-tuned on the SamSum dataset Gliwa et al. (2019), to output multi-sentence summaries of dialogue. For the high-level summarization module (Figure 1, row 5), we use BART, fine-tuned first for document summarization on the CNN/Daily Mail dataset Hermann et al. (2015), and then on SummScreen3D. The input for the latter fine-tuning is our re-ordered scene summaries and visual text descriptions, i.e., the output of the dialogue summarization, vision-processing and reordering modules. The output is the gold summaries from the SummScreen3D Papalampidi and Lapata (2023) training set.
@@ 5.2 Results
Table 1 summarizes our results, as measured by ROUGE-1 (r1), ROUGE-2 (r2), ROUGE-Lsum (rlsum) (computed using the python-rouge package) and our new PRisma metric. We present results with BERTScore Zhang et al. (2019) in Appendix E, for the sake of brevity. Results for Papalampidi and Lapata (2023), which we denote as ‘adapter-e2e’, were reproduced with their code. [^4] https://github.com/ppapalampidi/video_abstractive_summarization For PRisma, we use a separate query to GPT for each fact to check whether it is supported. There are roughly 70 facts per generated summary, which can lead to significant financial cost if used excessively. Therefore, while we report five random seeds for ROUGE, we select a single seed (randomly) for fact-precision and fact-recall.
@@
Here, in Table 3, we report an empirical test of the accuracy of our scene-detection algorithm (Section 3.2, Algorithm 1). These figures are produced using an episode which has explicitly marked scene breaks. We assign each transcript line a label based on the scene it is in with respect to these explicit scene breaks, e.g., all lines in the first scene get the label ‘0’. Then we do the same for the scene breaks produced by our algorithm, and compare the two sets of labels using unsupervised label comparison metrics, as commonly used in clustering problems: accuracy (ACC), normalized mutual information (NMI), and adjusted Rand index (ARI), defined as, e.g., in Mahon and Lukasiewicz (2023). We compare to two baselines. The first, denoted ‘uniform’, divides each episode into $n$ equally-sized scenes, where $n$ is the average number of scenes per episode. The second, denoted ‘uniform oracle’, does the same except sets $n$ to the true number of scenes in that episode.
@@
Humans prefer summaries generated by our modular approach
We selected 8 episodes from SummScreen3D, and asked human participants (recruited through the Prolific platform) to watch the episode, then rank summaries in order of perceived quality. The summaries we included in this study were ours, the gold summary, unlimiformer Bertsch et al. (2023), ‘adapter e2e‘ Papalampidi and Lapata (2023), and finetuned Mistral. We instructed participants to priortize informativeness and factual correctness, with a secondary emphasis on fluent well-written text. We included an attention-check question that participants would only be able to answer if they had watched the full episode, and asked for a short justification for the chosen ranking. The full instructions can be found in Appendix H. We recruited five annotators for each of the eight episodes.

### v9

@@ Abstract
In this paper we address the task of summarizing television shows, which touches key areas in AI research: complex reasoning, multiple modalities, and long narratives. We present a modular approach where separate components perform specialized sub-tasks which we argue affords greater flexibility compared to end-to-end methods. Our modules involve detecting scene boundaries, reordering scenes so as to minimize the number of cuts between different events, converting visual information to text, summarizing the dialogue in each scene, and fusing the scene summaries into a final summary for the entire episode. We also present a new metric, PRisma (Precision and Recall Evaluation of Summary Facts), to measure both precision and recall of generated summaries, which we decompose into atomic facts. Tested on the recently released SummScreen3D dataset Papalampidi and Lapata (2023), our method produces higher quality summaries than comparison models, as measured with ROUGE and our new fact-based metric, and as assessed by human evaluators. Code for our experiments and metric are at https://github.com/Lou1sM/modular_multimodal_summarization.
@@ 3.2 Scene Detection
We define a cost for a given partition, by invoking the minimum description length (MDL) principle Grünwald (2007), in which the optimal representation of a piece of data is that which contains the fewest of bits. MDL has been shown to have many nice theoretical properties (Grünwald, 2007) and has been used in recent machine learning applications from interpreting hidden representations Voita and Titov (2020) to measuring noise-free information content in images (Mahon and Lukasiewicz, 2024b; Mahon, 2024). We use it here to formalize the desirable properties of a good segmentation.
@@ 3.5 Summary Generation
Because of our two-layer summarization approach, we can use a relatively small backbone model, and are still able to encode very long input. In our experiments, we use variants of BART-large Lewis et al. (2020) for both the dialogue summarization and high-level summarization modules, but any other similar model could be used instead. For the dialogue summarization module (Figure 1, row 4), we use the public Huggingface checkpoint for BART which has been fine-tuned on the SamSum dataset Gliwa et al. (2019), to output multi-sentence summaries of dialogue. For the high-level summarization module (Figure 1, row 5), we use BART, fine-tuned first for document summarization on the CNN/Daily Mail dataset Hermann et al. (2015), and then on SummScreen3D. The input for the latter fine-tuning is our re-ordered scene summaries and visual text descriptions, i.e., the output of the dialogue summarization, vision-processing and reordering modules. The output is the gold summaries from the SummScreen3D training set.
@@ 5.2 Results
Table 1 summarizes our results, as measured by ROUGE-1 (r1), ROUGE-2 (r2), ROUGE-Lsum (rlsum) (computed using the python-rouge package) and our new PRisma metric. We present results with BERTScore Zhang et al. (2019) in Appendix E, for the sake of brevity. Results for Papalampidi and Lapata (2023), which we denote as ‘adapter-e2e’, were reproduced with their code. For PRisma, we use a separate query to GPT for each fact to check whether it is supported. There are roughly 70 facts per generated summary, which can lead to significant financial cost if used excessively. Therefore, while we report five random seeds for ROUGE, we select a single seed (randomly) for fact-precision and fact-recall.
@@
Here, in Table 3, we report an empirical test of the accuracy of our scene-detection algorithm (Section 3.2, Algorithm 1). These figures are produced using an episode which has explicitly marked scene breaks. We assign each transcript line a label based on the scene it is in with respect to these explicit scene breaks, e.g., all lines in the first scene get the label ‘0’. Then we do the same for the scene breaks produced by our algorithm, and compare the two sets of labels using unsupervised label comparison metrics, as commonly used in clustering problems Mahon and Lukasiewicz (2021): accuracy (ACC), normalized mutual information (NMI), and adjusted Rand index (ARI), defined as, e.g., in Mahon and Lukasiewicz (2024a). We compare to two baselines. The first, denoted ‘uniform’, divides each episode into $n$ equally-sized scenes, where $n$ is the average number of scenes per episode. The second, denoted ‘uniform oracle’, does the same except sets $n$ to the true number of scenes in that episode.
@@
Humans prefer summaries generated by our modular approach.
We selected 8 episodes from SummScreen3D, and asked human participants (recruited through the Prolific platform) to watch the episode, then rank the following summaries in order of perceived quality: ours, the gold summary, unlimiformer Bertsch et al. (2023), ‘adapter e2e‘ Papalampidi and Lapata (2023), and finetuned Mistral. We instructed participants to priortize informativeness and factual correctness, with a secondary emphasis on fluent well-written text. We included an attention-check question that participants would only be able to answer if they had watched the full episode, and asked for a short justification for the chosen ranking. The full instructions can be found in Appendix H. We recruited five annotators for each of the eight episodes.